# Wiman-Valiron method for fractional derivatives and sharp growth estimates of $\alpha$-analytic solutions for linear fractional differential equations

### v1

@@ 2.1 Auxiliary results from Wiman-Valiron theory
Lemma 2 (Lemma 3.4. [21], cf. Lemma 2 [12]).
@@
$\sum\limits_{|n-\nu|>\varkappa(\nu)}n^{q}|a_{n}|\rho^{n}=o\left(\frac{\nu^{q}\mu(r,\rho,f)}{v(\nu)^{3}}\right),\quad\nu=\nu(r,f),$ (1)
@@
Lemma 3 (Lemma 3.5. [21], cf. Lemma 7 [12]).
@@
Lemma 4 (Lemma 3.6. [21], cf. Lemma 8, [12]).
@@ 2.2 Chain rule for higher derivaitves and Bell polynomials
$(f\circ g)^{(n)}=\sum_{k=1}^{n}f^{(k)}(g)B_{nk}(g^{\prime},g^{\prime\prime},\dots,g^{(n-k+1)})$ (2)
@@
$B_{n,k}(z_{1},\dots,z_{n-k+1})=\sum_{\begin{subarray}{c}j_{1}+2j_{2}+\dots+(n-k+1)j_{n-k+1}=n\\ j_{1}+j_{2}+\dots+j_{n-k+1}=k\end{subarray}}\frac{n!}{j_{1}!\cdots j_{n-k+1}!}\prod_{s=1}^{n-k+1}\left(\frac{z_{s}}{s!}\right)^{j_{s}},$ (3)
@@
$B_{nk}^{*}(g^{\prime},g^{\prime\prime},\dots,g^{(n-k+1)})\leq(n-1)!\alpha^{k}(\alpha+1)\cdots(\alpha+n-k)|w|^{k\alpha-n}.$ (4)
@@
Assume that (4) holds for $n\leq m$ and $1\leq k\leq n$, and $k\leq m-1$. Then
@@
Combining this with (3) we deduce
@@
$+\sum_{\begin{subarray}{c}j_{1}+2j_{2}+\dots+(m-k+1)j_{m-k+1}=m\\ j_{1}+j_{2}+\dots+j_{m-k+1}=k\end{subarray}}\frac{m!}{j_{1}!\cdots j_{m-k+1}!}\sum_{s=1}^{m-k+1}j_{s}\frac{|g^{(s+1)}|}{|g^{(s)}|}\prod_{l=1}^{m-k+1}\left(\frac{|g^{(l)}|}{l!}\right)^{j_{l}}$ (5)
@@
We rewrite (5) as follows, using the induction assumption
@@
$=|g^{\prime}|B^{*}_{m,k-1}(g^{\prime},\dots,g^{(m-k+2)})+\frac{\alpha k+m}{|w|}B^{*}_{m,k}(g^{\prime},\dots,g^{(m-k+1)})$
@@
$=(m-1)!|w|^{k\alpha-m-1}\alpha^{k}(\alpha+1)\cdots(\alpha+m-k)\left[\alpha+m-k+1+\alpha k+m\right]$
@@ 3 Proof of Theorem 1
Since $\mu(r,\rho,f)\leq\mu(\rho,f)\leq M(\rho,f)$, from Lemma 2 with $q=0$ for all $\rho$, $|\ln\rho-\ln r|\leq\frac{1}{\varkappa(\nu)}$, we obtain
$f(z)=P(z)z^{\nu_{1}}+o\left(\frac{\mu(r,\rho,f)}{v(\nu)^{3}}\right)=P(z)z^{\nu_{1}}+o\left(\frac{M(\rho,f)}{v(\nu)^{3}}\right),\quad|z|=\rho$ (1)
@@
$P(z)=\sum\limits_{|n-\nu|\leq\varkappa(\nu)}|a_{n}|z^{n-\nu_{1}}.$ (2)
In particular, from (1) with $\rho=r$ we have $|P(z)|r^{\nu_{1}}\leq(1+o(1))M(r,f)$, i.e. for all sufficiently large $r\not\in E$
$|P(z)|\leq\frac{1,01M(r,f)}{r^{\nu_{1}}}=:M^{*}(r),\quad|z|=r.$ (3)
@@
$f(z)=P(z)z^{\nu_{1}}+R(z),$ (4)
where $P(z)$ is the polynomial (2). From now on we assume that $r$ is large enough so $\nu_{1}>\max\{j,\alpha j\}$.
Next, we need the asymptotic estimate of the Gamma functions ([24])
$\frac{\Gamma(t+a)}{\Gamma(t+b)}=t^{a-b}\left(1+O\left(\frac{1}{t}\right)\right),\quad t\rightarrow+\infty,\quad b,a\in\mathbb{R}.$ (5)
@@
$\leq C\sum\limits_{|n-\nu|>\varkappa(\nu)}(n\alpha)^{j\alpha}|a_{n}|\rho^{n-j}=o\left(\frac{\nu^{j\alpha}\mu(r,\rho,f)}{\rho^{j}v(\nu)^{3}}\right),\quad r\to\infty,r\not\in E.$ (6)
@@
$|P^{(q)}(z)|=O\left(\left(\frac{\varkappa(\nu)}{r}\right)^{q}\right)M^{*}(r).$ (7)
@@
$\leq\frac{2eM^{*}(r)\varkappa(\nu)}{\rho}\left(1+\frac{1}{40\varkappa(\nu)}\right)^{2\varkappa(\nu)}=O\left(\frac{\varkappa(\nu)}{r}M^{*}(r)\right).$
$|P^{(j)}(z)|\leq\frac{e\max\{|P^{(j-1)}(z)|\}(2\varkappa(\nu)-j)\rho^{\nu_{2}-\nu_{1}-j-1}}{r^{\nu_{2}-\nu_{1}-j}}=O\left(\left(\frac{\varkappa(\nu)}{r}\right)^{j}M^{*}(r)\right).$
We need the generalization Leibniz’s formula for fractional derivatives in order to estimate the fractional derivative of the first summand in (4). Let $f(x)$ and $g(x)$ be analytic functions on $[a,b]$. Then, according to ([20], p.216)
$D^{q}(f\cdot g)=\sum\limits_{k=0}^{+\infty}{q\choose k}(D^{q-k}f)g^{(k)},$ (8)
@@
$=Q((D^{\alpha})^{j}(w^{\alpha\nu_{1}}P(w^{\alpha})))=Q(D^{\alpha j}(w^{\alpha\nu_{1}}P(w^{\alpha}))).$ (9)
In the following arguments we consider a branch of the power function chosen on the segment $[0,w]$ emanating from the origin. Using (9) and (8) we obtain
@@
Applying Lemma 5 we get
@@
$=\frac{\alpha j\Gamma(\alpha\nu_{1}+1)|z|^{\nu_{1}-j}}{\Gamma(\alpha)|\Gamma(1-j\alpha)|}\sum_{k=1}^{2\varkappa(\nu)}|P^{(k)}(z)|\alpha^{k}|z|^{k}\sum_{m=k}^{\infty}\frac{|\Gamma(m-j\alpha)|\Gamma(\alpha+m-k+1)}{\Gamma(m+1)\Gamma(\alpha\nu_{1}+1-j\alpha+m)}.$ (10)
Let $b_{m}=\frac{|\Gamma(m-j\alpha)|\Gamma(\alpha+m-k+1)}{\Gamma(m+1)\Gamma(\alpha\nu_{1}+1-j\alpha+m)}$. To estimate the sum $\sum_{m=k}^{\infty}$ we consider two cases. First, let $k\leq m\leq[2\alpha\nu_{1}]$. Then
$\frac{b_{m+1}}{b_{m}}=\frac{|m-j\alpha|}{m+1}\frac{\alpha+m-k+1}{\alpha\nu_{1}+m+1-j\alpha}<\frac{2\alpha\nu_{1}+\alpha-k+1}{3\alpha\nu_{1}+1-j\alpha}<\frac{3}{4},\quad\nu\to\infty.$
So,
$\sum_{m=k}^{[2\alpha\nu_{1}]}b_{m}<4b_{k}=4\frac{|\Gamma(k-j\alpha)|\Gamma(\alpha+1)}{\Gamma(k+1)\Gamma(\alpha\nu_{1}+1-j\alpha+k)}.$ (11)
@@
$\Gamma(x)=x^{x-\frac{1}{2}}e^{-x}\sqrt{2\pi}e^{\frac{\theta(x)}{12x}},\quad\theta(x)\in(0,1),\;x\to+\infty,$ (12)
we deduce
@@
$\frac{\Gamma(\alpha+m-k+1)}{\Gamma(\alpha\nu_{1}+1-j\alpha+m)}\leq 2\frac{(\alpha\nu_{1}+m+1-j\alpha)^{\frac{1}{2}}}{{(\alpha+m-k+1)^{\alpha\nu_{1}+k-(j+1)\alpha+\frac{1}{2}}}}.$ (13)
@@
$\sum_{m=[2\alpha\nu_{1}]+1}^{\infty}b_{m}\leq\int_{2\alpha\nu_{1}}^{\infty}\left(\frac{2}{x}\right)^{\alpha\nu_{1}+k-\alpha+1}\,dx=\frac{2}{\alpha\nu_{1}+k-\alpha}\frac{1}{(\nu_{1}\alpha)^{\alpha\nu_{1}+k-\alpha}}.$
@@
Finally,
$\sum_{m=k}^{\infty}b_{m}\lesssim b_{k}=\frac{|\Gamma(k-j\alpha)|}{\Gamma(k+1)}\frac{\Gamma(\alpha+1)}{\Gamma(\alpha\nu_{1}+1+k-j\alpha)}.$ (14)
Taking into account (10), (7) and (5) we obtain
@@
$\lesssim|z|^{\nu_{1}-j}(\alpha\nu_{1})^{j\alpha}\sum\limits_{k=1}^{2\varkappa(\nu)}\left(\frac{\varkappa(\nu)\rho}{r\nu_{1}}\right)^{k}M^{*}(r)\lesssim|z|^{\nu_{1}-j}(\alpha\nu_{1})^{j\alpha}\frac{\varkappa(\nu)}{\nu}M^{*}(r)$
Therefore, in view of (1) and the previous estimate we have
@@
$=\frac{\Gamma(\alpha\nu_{1}+1)}{\Gamma(\alpha\nu_{1}+1-j\alpha)}\left(\frac{f(z)}{z^{j}}+o\left(\frac{\mu(r,\rho,f)}{\rho^{j}v(\nu)^{3}}\right)+O\left(\frac{\varkappa(\nu)}{\nu}\frac{M^{*}(r)}{\rho^{j}}\rho^{\nu_{1}}\right)\right).$ (15)
Since $\frac{1}{v(t)^{3}}=o\left(\frac{\varkappa(t)}{t}\right)$, $t\rightarrow+\infty$, using (15) and (6) we obtain for $|z|=\rho$
@@
$=\frac{\Gamma(\nu_{1}\alpha+1)}{\Gamma(\nu_{1}\alpha+1-j\alpha)}\left(\frac{f(z)}{z^{j}}+o\left(\frac{\varkappa(\nu)}{\nu}\frac{M(\rho,f)}{\rho^{j}}\right)+O\left(\frac{\varkappa(\nu)}{\nu}\frac{M(r,f)}{\rho^{j}}\left(\frac{\rho}{r}\right)^{\nu_{1}}\right)\right)$ (16)
@@
Next we choose $z_{0}$ so that $|f(z_{0})|=M(r,f)$ and take $\eta=1$, $\tau=\ln(\rho/r)$. Then, by Lemma B, we have
@@
Thus, (16) yields
$\mathcal{D}^{j}_{\alpha}f(z)=\frac{\Gamma(\nu_{1}\alpha+1)}{\Gamma(\nu_{1}\alpha+1-j\alpha)}\left(\frac{f(z)}{z^{j}}+O\left(\frac{\varkappa(\nu)}{\nu}\frac{M(\rho,f)}{\rho^{j}}\right)\right).$ (17)
From (5) we have
$\frac{\Gamma(\nu_{1}\alpha+1)}{\Gamma(\nu_{1}\alpha+1-j\alpha)}=(\nu\alpha)^{j\alpha}\left(1+O\left(\frac{1}{\nu}\right)\right),\quad\nu\rightarrow+\infty.$ (18)
Therefore, (18) implies
@@
when $r\rightarrow+\infty$ outside a set of finite logarithmic measure, which is (12).
We then choose $z$ in (12) to maximise $|f(z)|$ and $|\mathcal{D}^{\alpha}f(z)|$ and deduce that
$M(\rho,\mathcal{D}^{j}_{\alpha}f)=\left(1+O\left(\frac{\varkappa(\nu)}{\nu}\right)\right)\frac{(\nu\alpha)^{j\alpha}}{\rho^{j}}M(\rho,f).$
To complete the proof of (13) it remains to show that
$\ln M(\rho,f)=\ln M(r,f)+\nu\ln(\rho/r)+o(1).$
To see this we note that (4) and (6) yield for our range of $\rho$
$\ln M(\rho,f)=\nu_{1}\ln\rho+\ln M(\rho,P)+o(1).$
On the other hand it follows from Lemma 3 that
$M(\rho,P)=M(r,P)\left(1+O\left(\frac{(\rho-r)\varkappa(\nu)}{r}\right)\right)\sim M(r,P)$
if $\varkappa(\nu)\ln(\rho/r)=o(1)$. The second equality of (13) now follows, completing the proof of Theorem 1.
4 $\alpha$-analyticity of soultions for (1)
Using the Cauchy method of majorant series, we prove the existence and uniqueness of a solution to (1).
@@ 4 $\alpha$-analyticity of soultions for (1)
The equation (1) where $p_{k}(x)=\sum_{m=0}^{\infty}p_{mk}x^{m\alpha}$, $x\in[0,\rho_{k})$ are $\alpha$-analytic, $k\in\{0,\dots,n-1\}$ with the initial conditions
$y(0)=b_{0},\;\mathbb{D}_{\alpha}y(0)=b_{1},\dots,\mathbb{D}_{\alpha}^{n-1}y(0)=b_{n-1},$ (1)
@@
$\mathbb{D}_{\alpha}^{k}y(x)=\sum_{m=0}^{\infty}a_{m+k}\frac{\Gamma((m+k)\alpha+1)}{\Gamma(m\alpha+1)}x^{m\alpha}.$ (2)
This yields $\mathbb{D}_{\alpha}^{k}y(0)=a_{k}\Gamma(k\alpha+1)$. Hence, $a_{k}=b_{k}/\Gamma(k\alpha+1)$, $k\in\{0,\dots,n-1\}$. Substituting (2) into (1) we obatin
@@
$=-\sum_{k=0}^{n-1}x^{m\alpha}\sum_{s=0}^{m}a_{s+k}\frac{\Gamma((s+k)\alpha+1)}{\Gamma(s\alpha+1)}p_{m-s,k}.$ (3)
Equating the coefficients of the same degree in (3), we write
$a_{m+n}\frac{\Gamma((m+n)\alpha+1)}{\Gamma(m\alpha+1)}=-\sum_{k=0}^{n-1}\sum_{s=0}^{m}a_{s+k}p_{m-s,k}\frac{\Gamma((s+k)\alpha+1)}{\Gamma(s\alpha+1)},\quad m\in\mathbb{Z}_{+}.$ (4)
@@
$|p_{j,k}|\leq\frac{M}{r^{j\alpha}},\quad j\in\mathbb{Z}_{+},\;k\in\{0,1,\dots,n-1\}.$ (5)
@@
$|a_{p}|\leq\beta_{p}\prod_{j=1}^{p-1}\left(\frac{1}{r^{\alpha}}+\beta_{j}\right)$ (6)
@@
It follows from properties of the Gamma-function that
$\frac{\Gamma(m\alpha+1)}{\Gamma((m+n)\alpha+1)}\leq\frac{M_{1}}{(m\alpha+1)^{n\alpha}}=:\gamma_{m},\quad m\in\mathbb{Z}_{+},n\in\mathbb{N}$ (7)

### v2

@@ 2.1 Auxiliary results from Wiman-Valiron theory
Lemma 2 ([21, Lemma 3.4], cf. [12, Lemma 2]).
@@
$\sum\limits_{|n-\nu|>\varkappa(\nu)}n^{q}|a_{n}|\rho^{n}=o\left(\frac{\nu^{q}\mu(r,\rho,f)}{v(\nu)^{3}}\right),\quad\nu=\nu(r,f),$ (2.1)
@@
Lemma 3 ([21, Lemma 3.5], cf. [12, Lemma 7]).
@@
Lemma 4 ([21, Lemma 3.6], cf. [12, Lemma 8]).
@@ 2.2 Chain rule for higher derivaitves and Bell polynomials
$(f\circ g)^{(n)}=\sum_{k=1}^{n}f^{(k)}(g)B_{n,k}(g^{\prime},g^{\prime\prime},\dots,g^{(n-k+1)})$ (2.2)
@@
$B_{n,k}(z_{1},\dots,z_{n-k+1})=\sum_{\begin{subarray}{c}j_{1}+2j_{2}+\dots+(n-k+1)j_{n-k+1}=n\\ j_{1}+j_{2}+\dots+j_{n-k+1}=k\end{subarray}}\frac{n!}{j_{1}!\cdots j_{n-k+1}!}\prod_{s=1}^{n-k+1}\left(\frac{z_{s}}{s!}\right)^{j_{s}},$ (2.3)
@@
$B_{n,k}^{*}(g^{\prime},g^{\prime\prime},\dots,g^{(n-k+1)})\leq(n-1)!\alpha^{k}(\alpha+1)\cdots(\alpha+n-k)|w|^{k\alpha-n}.$ (2.4)
@@
Assume that (2.4) holds for $n\leq m$ and $1\leq k\leq n$, and $k\leq m-1$. The case $k=m=n$ is already considered. Then
@@
Combining this with (2.3) we deduce
@@
$+\sum_{\begin{subarray}{c}j_{1}+2j_{2}+\dots+(m-k+1)j_{m-k+1}=m\\ j_{1}+j_{2}+\dots+j_{m-k+1}=k\end{subarray}}\frac{m!}{j_{1}!\cdots j_{m-k+1}!}\sum_{s=1}^{m-k+1}j_{s}\frac{|g^{(s+1)}|}{|g^{(s)}|}\prod_{l=1}^{m-k+1}\left(\frac{|g^{(l)}|}{l!}\right)^{j_{l}}$ (2.5)
@@
We rewrite (2.5) as follows, using the induction assumption,
@@
$\leq|g^{\prime}|B^{*}_{m,k-1}(g^{\prime},\dots,g^{(m-k+2)})+\frac{\alpha k+m}{|w|}B^{*}_{m,k}(g^{\prime},\dots,g^{(m-k+1)})$
@@
$=(m-1)!|w|^{k\alpha-m-1}\alpha^{k}(\alpha+1)\cdots(\alpha+m-k)\left(\alpha+m-k+1+\alpha k+m\right)$
@@ 3 Proof of Theorem 1
Since, by Cauchy’s inequality, $\mu(r,\rho,f)\leq\mu(\rho,f)\leq M(\rho,f)$, from Lemma 2 with $q=0$ for all $\rho$, $|\ln\rho-\ln r|\leq\frac{1}{\varkappa(\nu)}$, we obtain
$f(z)=P(z)z^{\nu_{1}}+o\left(\frac{\mu(r,\rho,f)}{v(\nu)^{3}}\right)=P(z)z^{\nu_{1}}+o\left(\frac{M(\rho,f)}{v(\nu)^{3}}\right),\quad|z|=\rho$ (3.1)
@@
$P(z)=\sum\limits_{|n-\nu|\leq\varkappa(\nu)}|a_{n}|z^{n-\nu_{1}}.$ (3.2)
From (3.1) with $\rho=r$ we have $|P(z)|r^{\nu_{1}}\leq(1+o(1))M(r,f)$, $r\to\infty$, $r\not\in E$, i.e. for all sufficiently large $r\not\in E$
$|P(z)|\leq\frac{1,01M(r,f)}{r^{\nu_{1}}}=:M^{*}(r),\quad|z|=r.$ (3.3)
@@
$f(z)=P(z)z^{\nu_{1}}+R(z),$ (3.4)
where $P(z)$ is the polynomial (3.2). From now on we assume that $r$ is large enough so $\nu_{1}>\max\{j,\alpha j\}$.
Next, we need the asymptotic estimate of the Gamma function ([24])
$\frac{\Gamma(t+a)}{\Gamma(t+b)}=t^{a-b}\left(1+O\left(\frac{1}{t}\right)\right),\quad t\rightarrow+\infty,\quad b,a\in\mathbb{R}.$ (3.5)
@@
$\leq C\sum\limits_{|n-\nu|>\varkappa(\nu)}(n\alpha)^{j\alpha}|a_{n}|\rho^{n-j}=o\left(\frac{\nu^{j\alpha}\mu(r,\rho,f)}{\rho^{j}v(\nu)^{3}}\right),\quad r\to\infty,r\not\in E.$ (3.6)
@@
$|P^{(q)}(z)|=O\left(\left(\frac{\varkappa(\nu)}{r}\right)^{q}M^{*}(r)\right).$ (3.7)
@@
$\leq\frac{2eM^{*}(r)\varkappa(\nu)}{\rho}\left(1+\frac{1}{40\varkappa(\nu)}\right)^{2\varkappa(\nu)}=O\left(\frac{\varkappa(\nu)}{r}M^{*}(r)\right),\quad r\to\infty.$
Then
$|P^{(j)}(z)|\leq\frac{eM(\rho,P^{(j-1)})(2\varkappa(\nu)-j)\rho^{\nu_{2}-\nu_{1}-j-1}}{r^{\nu_{2}-\nu_{1}-j}}=O\left(\left(\frac{\varkappa(\nu)}{r}\right)^{j}M^{*}(r)\right).$
We need the generalization Leibniz’s formula for fractional derivatives in order to estimate the fractional derivative of the first summand in (3.4). Let $f(x)$ and $g(x)$ be analytic functions on $[a,b]$. Then, according to ([20, p. 216]),
$D^{q}(f\cdot g)=\sum\limits_{k=0}^{+\infty}{q\choose k}(D^{q-k}f)g^{(k)},$ (3.8)
@@
$=Q((D^{\alpha})^{j}(w^{\alpha\nu_{1}}P(w^{\alpha})))=Q(D^{\alpha j}(w^{\alpha\nu_{1}}P(w^{\alpha}))),\quad|z|\to\infty.$ (3.9)
In the following arguments we consider a branch of the power function chosen on the segment $[0,w]$ emanating from the origin. Using (3.9) and (3.8) we obtain
@@
Applying Lemma 5 and recalling that $\deg P\leq 2\varkappa(\nu)$ we get
@@
$=\frac{\alpha j\Gamma(\alpha\nu_{1}+1)|z|^{\nu_{1}-j}}{\Gamma(\alpha)|\Gamma(1-j\alpha)|}\sum_{k=1}^{2\varkappa(\nu)}|P^{(k)}(z)|\alpha^{k}|z|^{k}\sum_{m=k}^{\infty}\frac{|\Gamma(m-j\alpha)|\Gamma(\alpha+m-k+1)}{\Gamma(m+1)\Gamma(\alpha\nu_{1}+1-j\alpha+m)}.$ (3.10)
Let $b_{m}=\frac{|\Gamma(m-j\alpha)|\Gamma(\alpha+m-k+1)}{\Gamma(m+1)\Gamma(\alpha\nu_{1}+1-j\alpha+m)}$. To estimate the sum $\sum_{m=k}^{\infty}$ we consider two cases. First, let $k\leq m\leq[2\alpha\nu_{1}]$. If $m\geq j\alpha$, then
$\frac{b_{m+1}}{b_{m}}=\frac{m-j\alpha}{m+1}\frac{\alpha+m-k+1}{\alpha\nu_{1}+m+1-j\alpha}<\frac{2\alpha\nu_{1}+\alpha-k+1}{3\alpha\nu_{1}+1-j\alpha}<\frac{3}{4},\quad\nu\to\infty.$
Otherwise,
$\frac{b_{m+1}}{b_{m}}\leq\frac{j\alpha}{2}\frac{\alpha+j\alpha+1}{\alpha\nu_{1}+2-j\alpha},$
and we arrive to the same conclusion as $r\to\infty$ because $\nu_{1}\to\infty$. So,
$\sum_{m=k}^{[2\alpha\nu_{1}]}b_{m}<4b_{k}=4\frac{|\Gamma(k-j\alpha)|\Gamma(\alpha+1)}{\Gamma(k+1)\Gamma(\alpha\nu_{1}+1-j\alpha+k)}.$ (3.11)
@@
$\Gamma(x)=x^{x-\frac{1}{2}}e^{-x}\sqrt{2\pi}e^{\frac{\theta(x)}{12x}},\quad\theta(x)\in(0,1),\;x\to+\infty,$ (3.12)
we deduce $(\theta_{j}(x)\in(0,1),j\in\{1,2\})$
@@
$\frac{\Gamma(\alpha+m-k+1)}{\Gamma(\alpha\nu_{1}+1-j\alpha+m)}\leq 2\frac{(\alpha\nu_{1}+m+1-j\alpha)^{\frac{1}{2}}}{{(\alpha+m-k+1)^{\alpha\nu_{1}+k-(j+1)\alpha+\frac{1}{2}}}}.$ (3.13)
@@
$\sum_{m=[2\alpha\nu_{1}]+1}^{\infty}b_{m}\leq\int_{2\alpha\nu_{1}}^{\infty}\left(\frac{2}{x}\right)^{\alpha\nu_{1}+k-\alpha+1}\,dx=\frac{2}{\alpha\nu_{1}+k-\alpha}\frac{1}{(\nu_{1}\alpha)^{\alpha\nu_{1}+k-\alpha}}.$ (3.14)
@@
Combining this with (3.11) and (3.14), we obtain
$\sum_{m=k}^{\infty}b_{m}\lesssim b_{k}=\frac{|\Gamma(k-j\alpha)|}{\Gamma(k+1)}\frac{\Gamma(\alpha+1)}{\Gamma(\alpha\nu_{1}+1+k-j\alpha)}.$ (3.15)
Taking into account (3.10), (3.7) and (3.5) we obtain
@@
$\lesssim|z|^{\nu_{1}-j}(\alpha\nu_{1})^{j\alpha}\sum\limits_{k=1}^{2\varkappa(\nu)}\left(\frac{\varkappa(\nu)\rho}{r\nu_{1}}\right)^{k}M^{*}(r)\lesssim|z|^{\nu_{1}-j}(\alpha\nu_{1})^{j\alpha}\frac{\varkappa(\nu)}{\nu}M^{*}(r).$
Therefore, in view of (3.1) and the previous estimate we have
@@
$=\frac{\Gamma(\alpha\nu_{1}+1)}{\Gamma(\alpha\nu_{1}+1-j\alpha)}\left(\frac{f(z)}{z^{j}}+o\left(\frac{\mu(r,\rho,f)}{\rho^{j}v(\nu)^{3}}\right)+O\left(\frac{\varkappa(\nu)}{\nu}\frac{M^{*}(r)}{\rho^{j}}\rho^{\nu_{1}}\right)\right).$ (3.16)
Since $\int_{0}^{\infty}\frac{dt}{v(t)}<\infty$ and $v$ is nondecreasing, $v(t)/t\to+\infty$ as $t\to+\infty$. Hence $\frac{1}{v(t)^{3}}=o\left(\frac{\varkappa(t)}{t}\right)$, $t\rightarrow+\infty$, and using (3.16) and (3.6) we obtain for $|z|=\rho$
@@
$=\frac{\Gamma(\nu_{1}\alpha+1)}{\Gamma(\nu_{1}\alpha+1-j\alpha)}\left(\frac{f(z)}{z^{j}}+o\left(\frac{\varkappa(\nu)}{\nu}\frac{M(\rho,f)}{\rho^{j}}\right)+O\left(\frac{\varkappa(\nu)}{\nu}\frac{M(r,f)}{\rho^{j}}\left(\frac{\rho}{r}\right)^{\nu_{1}}\right)\right)$ (3.17)
@@
Next we choose $z_{0}$ so that $|f(z_{0})|=M(r,f)$ and take $\tau=\ln(\rho/r)$, $\eta=1$. Then, by Theorem B, we have
@@
Thus, (3.17) yields
$\mathcal{D}^{j}_{\alpha}f(z)=\frac{\Gamma(\nu_{1}\alpha+1)}{\Gamma(\nu_{1}\alpha+1-j\alpha)}\left(\frac{f(z)}{z^{j}}+O\left(\frac{\varkappa(\nu)}{\nu}\frac{M(\rho,f)}{\rho^{j}}\right)\right).$ (3.18)
From (3.5) we have
$\frac{\Gamma(\nu_{1}\alpha+1)}{\Gamma(\nu_{1}\alpha+1-j\alpha)}=(\nu\alpha)^{j\alpha}\left(1+O\left(\frac{1}{\nu}\right)\right),\quad\nu\rightarrow+\infty.$ (3.19)
Therefore, (3.19) implies
@@
when $r\rightarrow+\infty$ outside a set of finite logarithmic measure, which is (1.12).
We then choose $z$ in (1.12) in turn to maximise $|f(z)|$ and $|\mathcal{D}^{\alpha}f(z)|$ and deduce that
$M(\rho,\mathcal{D}^{j}_{\alpha}f)=\left(1+O\left(\frac{\varkappa(\nu)}{\nu}\right)\right)\frac{(\nu\alpha)^{j\alpha}}{\rho^{j}}M(\rho,f),\quad\rho\to+\infty.$
In order to complete the proof of (1.13) it is sufficient to show that
$\ln M(\rho,f)=\ln M(r,f)+\nu\ln(\rho/r)+o(1),\quad r\to+\infty.$
First, we note that (3.4) and (3.6) yield for our range of $\rho$
$\ln M(\rho,f)=\nu_{1}\ln\rho+\ln M(\rho,P)+o(1),\quad\rho\to+\infty.$
Then, it follows from Lemma 3 that
$M(\rho,P)=M(r,P)\left(1+O\left(\frac{(\rho-r)\varkappa(\nu)}{r}\right)\right)\sim M(r,P),\quad r\to+\infty.$
if $\varkappa(\nu)\ln(\rho/r)=o(1)$. The second equality of (1.13) now follows, completing the proof of Theorem 1.
4 $\alpha$-analyticity of soultions for (1.1)
Using the Cauchy method of majorant series, we prove the existence and uniqueness of a solution to (1.1).
@@
The equation (1.1) where $p_{k}(x)=\sum_{m=0}^{\infty}p_{mk}x^{m\alpha}$, $x\in[0,\rho_{k})$ are $\alpha$-analytic, $k\in\{0,\dots,n-1\}$ with the initial conditions
$y(0)=b_{0},\;\mathbb{D}_{\alpha}y(0)=b_{1},\dots,\mathbb{D}_{\alpha}^{n-1}y(0)=b_{n-1},$ (4.1)
@@
$\mathbb{D}_{\alpha}^{k}y(x)=\sum_{m=0}^{\infty}a_{m+k}\frac{\Gamma((m+k)\alpha+1)}{\Gamma(m\alpha+1)}x^{m\alpha}.$ (4.2)
This yields $\mathbb{D}_{\alpha}^{k}y(0)=a_{k}\Gamma(k\alpha+1)$. Hence, $a_{k}=b_{k}/\Gamma(k\alpha+1)$, $k\in\{0,\dots,n-1\}$. Substituting (4.2) into (1.1) we obtain
@@
$=-\sum_{k=0}^{n-1}x^{m\alpha}\sum_{s=0}^{m}a_{s+k}\frac{\Gamma((s+k)\alpha+1)}{\Gamma(s\alpha+1)}p_{m-s,k}.$ (4.3)
Equating the coefficients of the same degree in (4.3), we write
$a_{m+n}\frac{\Gamma((m+n)\alpha+1)}{\Gamma(m\alpha+1)}=-\sum_{k=0}^{n-1}\sum_{s=0}^{m}a_{s+k}p_{m-s,k}\frac{\Gamma((s+k)\alpha+1)}{\Gamma(s\alpha+1)},\quad m\in\mathbb{Z}_{+}.$ (4.4)
@@
$|p_{j,k}|\leq\frac{M}{r^{j\alpha}},\quad j\in\mathbb{Z}_{+},\;k\in\{0,1,\dots,n-1\}.$ (4.5)
@@
$|a_{p}|\leq\beta_{p}\prod_{j=1}^{p-1}\left(\frac{1}{r^{\alpha}}+\beta_{j}\right)$ (4.6)
@@
It follows from properties of the Gamma function that
$\frac{\Gamma(m\alpha+1)}{\Gamma((m+n)\alpha+1)}\leq\frac{M_{1}}{(m\alpha+1)^{n\alpha}}=:\gamma_{m},\quad m\in\mathbb{Z}_{+},n\in\mathbb{N}$ (4.7)